OpenStudy (anonymous):

1. Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. What would be the function of this? And What would be a reasonable domain for this function based on this scenario? Explain why this is a proper domain, and why this is the function.

OpenStudy (imstuck):

The function of this would be f(c) = 85 + 15x, which means in words, the function in terms of cost is equal to 85 dollars for the bracelet alone plus 15 dollars for each charm (which I designated as x. X is the number of charms.)

OpenStudy (anonymous):

Where it asks what the function it, it should say function notation.

OpenStudy (imstuck):

See the main cost for the bracelet is 85, no matter if you put 0 charms on it or 100 charms on it. What you spend on the bracelet depends upon how many charms you buy. That is in function notation. Unless you want me to use f(x) = 85 + 15x. That look better?

OpenStudy (imstuck):

The first question, "What would be the function of this?" That answer is f(x) = 85 + 15x.

OpenStudy (anonymous):

Okay i had f(x)=15x+85

OpenStudy (imstuck):

Domain defines x values. Just like the range defines y values. Here, when they ask for a reasonable domain, they are asking for what is a reasonable x value.

OpenStudy (imstuck):

15x + 85 is the same as 85 + 15x. As long as you have the x with the 15!

OpenStudy (anonymous):

so basically i just add 85 and 15x everytime to get the domain?

OpenStudy (imstuck):

I'M not sure how they want the domain noted. If it is in an inequality statement, it would be reasonable for \[x \ge85\]I say that because even without the charms, the trio will spend at least $85. I'm not sure how they want it noted. Any ideas from any other problems like this in classs?

OpenStudy (imstuck):

In fact, use that for the domain. The x is greater than or equal to 85. That's a reasonable domain.

OpenStudy (imstuck):

This is a reasonable domain because, like I said, no matter if they buy any charms or not, they will spend at least $85.

OpenStudy (imstuck):

As for "Why is this the function?" The function is adding the base bracelet of $85 to each charm, x, that cost $15 apiece.

OpenStudy (imstuck):

That answers all the questions in your problem. Do you understand?